What is the standard enthalpy change for vaporizing 4.000 mol C5H12O(L) at 25°C?

ΔH° = ________ kJ
1
Expert's answer
2019-04-02T05:57:34-0400
"C_5H_{12}O(l) \rightarrow C_5H_{12}O(g)"

"\Delta H_f^0(C_5H_{12}O(l)) = -351.6 \frac{kJ}{mol}"


"\Delta H_f^0 (C_5H_{12}O(g)) = -294.6 \frac{kJ}{mol}"



"\Delta H_{vap} = \sum \Delta H_{products} - \sum \Delta H_{reactants}"

"\Delta H_{vap} = \Delta H_f^0 (C_5H_{12}O(g))-\Delta H_f^0(C_5H_{12}O(l))= -294.6 \frac{kJ}{mol} -(-351.6 \frac{kJ}{mol}) = 57 \frac{kJ}{mol}"

"\Delta H^0 = \Delta H_{vap} \times n = 57\frac{kJ}{mol}\times 4.000 mol = 228.0 kJ"